The radius of the moon is one fourth and its mass is one eight first that of the earth, if the acceleration due to gravity on the surface of the earth is 9.8, what is the value on the moon surface.
1
Expert's answer
2019-10-10T09:42:37-0400

The acceleration due to gravity on the moon surface:


"g'=G\frac{M'}{R'^2}=G\frac{\frac{1}{81}M}{(\frac{1}{4}R)^2}=\frac{16}{81}G\frac{M}{R^2}"

"g'=\frac{16}{81}g=\frac{16}{81}9.8=1.9\frac{m}{s^2}"